Consider the IVP y'' + 3y' + 3y = (1 − u(t − 4)) with y'(0) = 0 and y(0) = 0. Solve the differential equation, and if possible, provide a graph
Answer #1

4 +34 +34 - 1 - ult-4); 460) = 460) = 0 Kucking Laplace transform Lly) + 3 Aly] + 3 fly] = e() - sculta) 2 szkly) - s 90 -y>Lly) - - (5+3/2) (5+3/2) 5+2+3/4 Siking Inverse Laplace Transform 7 g = Faces 3) Mey:-)! *****8.74(89 + L 1294 = * (94) La Sa (Git% bováha, **(31» (ová) + (-1) (5-7 > 349-cay- Salut) yay 2 Pas(fat 43 * ,24 9 sin (Oct-2) - Yuth) 909-